# Model Lakes

### v1

@@ 7. Conclusion
The database community has responded to the “Big Model” revolution by proposing platforms like Agora (Traub et al., 2020), that manage data-related assets, including models, datasets, software, and compute resources in a coherent ecosystem. There is now a need for general methods for managing model lakes. It will be important that the methods generalize, irrespective of how many models we are trying to understand, how they are trained, or what base of information we wish to use to resolve our tasks. We call on the database community to contribute to the vision of model lakes, supporting users to better distinguish a set of models that they might want to choose from for their own tasks as well as alleviate future data and model management concerns that come up the rise of deep learning models. Our vision is for a fundamentally new platform that extends and integrates works on data/model provenance, data/model citation, and data/model version management.
@@ Acknowledgements.
KP and RM was supported in part by NSF award numbers IIS-2107248, IIS-1956096, IIS-2325632, and KP and DB by a grant from Open Philanthropy.References
@@
- Alrashed et al. (2021) Tarfah Alrashed, Dimitris Paparas, Omar Benjelloun, Ying Sheng, and Natasha F. Noy. 2021. Dataset or Not? A Study on the Veracity of Semantic Markup for Dataset Pages. In The Semantic Web - ISWC 2021 - 20th International Semantic Web Conference, ISWC 2021, Virtual Event, October 24-28, 2021, Proceedings (Lecture Notes in Computer Science), Andreas Hotho, Eva Blomqvist, Stefan Dietze, Achille Fokoue, Ying Ding, Payam M. Barnaghi, Armin Haller, Mauro Dragoni, and Harith Alani (Eds.), Vol. 12922. Springer, 338–356. https://doi.org/10.1007/978-3-030-88361-4_20
@@
- Belrose et al. (2023) Nora Belrose, David Schneider-Joseph, Shauli Ravfogel, Ryan Cotterell, Edward Raff, and Stella Biderman. 2023. LEACE: Perfect linear concept erasure in closed form. arXiv:2306.03819 [cs.LG]
- Benjelloun et al. (2020) Omar Benjelloun, Shiyu Chen, and Natasha F. Noy. 2020. Google Dataset Search by the Numbers. In The Semantic Web - ISWC 2020 - 19th International Semantic Web Conference, Athens, Greece, November 2-6, 2020, Proceedings, Part II (Lecture Notes in Computer Science), Jeff Z. Pan, Valentina A. M. Tamma, Claudia d’Amato, Krzysztof Janowicz, Bo Fu, Axel Polleres, Oshani Seneviratne, and Lalana Kagal (Eds.), Vol. 12507. Springer, 667–682. https://doi.org/10.1007/978-3-030-62466-8_41
- Bernstein et al. (2000) Philip A. Bernstein, Alon Y. Halevy, and Rachel Pottinger. 2000. A Vision of Management of Complex Models. SIGMOD Rec. 29, 4 (2000), 55–63. https://doi.org/10.1145/369275.369289
- Bhardwaj et al. (2015) Anant P. Bhardwaj, Amol Deshpande, Aaron J. Elmore, David R. Karger, Sam Madden, Aditya G. Parameswaran, Harihar Subramanyam, Eugene Wu, and Rebecca Zhang. 2015. Collaborative Data Analytics with DataHub. Proc. VLDB Endow. 8, 12 (2015), 1916–1919. https://doi.org/10.14778/2824032.2824100
- Borisov et al. (2022) Vadim Borisov, Tobias Leemann, Kathrin Seßler, Johannes Haug, Martin Pawelczyk, and Gjergji Kasneci. 2022. Deep Neural Networks and Tabular Data: A Survey. IEEE Transactions on Neural Networks and Learning Systems (2022), 1–21. https://doi.org/10.1109/TNNLS.2022.3229161
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems 33 (2020), 1877–1901.
@@
- Casper et al. (2024) Stephen Casper, Carson Ezell, Charlotte Siegmann, Noam Kolt, Taylor Lynn Curtis, Benjamin Bucknall, Andreas Haupt, Kevin Wei, Jérémy Scheurer, Marius Hobbhahn, Lee Sharkey, Satyapriya Krishna, Marvin Von Hagen, Silas Alberti, Alan Chan, Qinyi Sun, Michael Gerovitch, David Bau, Max Tegmark, David Krueger, and Dylan Hadfield-Menell. 2024. Black-Box Access is Insufficient for Rigorous AI Audits. arXiv:2401.14446 [cs.CY]
- Chapman et al. (2020) Adriane Chapman, Elena Simperl, Laura Koesten, George Konstantinidis, Luis-Daniel Ibáñez, Emilia Kacprzak, and Paul Groth. 2020. Dataset search: a survey. VLDB J. 29, 1 (2020), 251–272. https://doi.org/10.1007/S00778-019-00564-X
@@
- Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Alex Castro-Ros, Marie Pellat, Kevin Robinson, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Yu, Vincent Zhao, Yanping Huang, Andrew Dai, Hongkun Yu, Slav Petrov, Ed H. Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V. Le, and Jason Wei. 2022. Scaling Instruction-Finetuned Language Models. arXiv:2210.11416 [cs.LG]
@@
- Conmy et al. (2023) Arthur Conmy, Augustine N. Mavor-Parker, Aengus Lynch, Stefan Heimersheim, and Adrià Garriga-Alonso. 2023. Towards Automated Circuit Discovery for Mechanistic Interpretability. arXiv:2304.14997 [cs.LG]
- Cun et al. (1990) Y. Le Cun, B. Boser, J. S. Denker, R. E. Howard, W. Habbard, L. D. Jackel, and D. Henderson. 1990. Handwritten digit recognition with a back-propagation network. Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 396–404.
- Dai et al. (2022) Damai Dai, Li Dong, Yaru Hao, Zhifang Sui, Baobao Chang, and Furu Wei. 2022. Knowledge Neurons in Pretrained Transformers. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Dublin, Ireland, 8493–8502. https://doi.org/10.18653/v1/2022.acl-long.581
- De Cao et al. (2021) Nicola De Cao, Wilker Aziz, and Ivan Titov. 2021. Editing Factual Knowledge in Language Models. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 6491–6506. https://doi.org/10.18653/v1/2021.emnlp-main.522
@@
- Dibia (2023) Victor Dibia. 2023. LIDA: A Tool for Automatic Generation of Grammar-Agnostic Visualizations and Infographics using Large Language Models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics: System Demonstrations, ACL 2023, Toronto, Canada, July 10-12, 2023, Danushka Bollegala, Ruihong Huang, and Alan Ritter (Eds.). Association for Computational Linguistics, 113–126. https://doi.org/10.18653/V1/2023.ACL-DEMO.11
- Doan et al. (2012) AnHai Doan, Alon Y. Halevy, and Zachary G. Ives. 2012. Principles of Data Integration. Morgan Kaufmann. http://research.cs.wisc.edu/dibook/
@@
- Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. 2023. Who’s Harry Potter? Approximate Unlearning in LLMs. arXiv:2310.02238 [cs.CL]
@@
- Fan et al. (2023a) Grace Fan, Jin Wang, Yuliang Li, and Renée J. Miller. 2023a. Table Discovery in Data Lakes: State-of-the-art and Future Directions. In Companion of the 2023 International Conference on Management of Data (Seattle, WA, USA) (SIGMOD ’23). Association for Computing Machinery, New York, NY, USA, 69–75. https://doi.org/10.1145/3555041.3589409
@@
- Ha et al. (2017) David Ha, Andrew M. Dai, and Quoc V. Le. 2017. HyperNetworks. In International Conference on Learning Representations. https://openreview.net/forum?id=rkpACe1lx
- Halevy et al. (2023) Alon Y. Halevy, Yejin Choi, Avrilia Floratou, Michael J. Franklin, Natasha F. Noy, and Haixun Wang. 2023. Will LLMs reshape, supercharge, or kill data science? Proc. VLDB Endow. 16, 12 (2023), 4114–4115. https://doi.org/10.14778/3611540.3611634
@@
- Kingma and Ba (2015) Diederik P. Kingma and Jimmy Ba. 2015. Adam: A Method for Stochastic Optimization. In 3rd International Conference on Learning Representations, ICLR 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track Proceedings, Yoshua Bengio and Yann LeCun (Eds.). http://arxiv.org/abs/1412.6980
@@
- Labs ([n.d.]a) Neptune Labs. [n.d.]a. neptune.ai — The MLOps stack component for experiment tracking. https://neptune.ai/.
@@
- Li et al. (2020) Yuliang Li, Jinfeng Li, Yoshihiko Suhara, AnHai Doan, and Wang-Chiew Tan. 2020. Deep Entity Matching with Pre-Trained Language Models. 14, 1 (sep 2020), 50–60. https://doi.org/10.14778/3421424.3421431
@@
- Liang et al. (2024) Weixin Liang, Nazneen Rajani, Xinyu Yang, Ezinwanne Ozoani, Eric Wu, Yiqun Chen, Daniel Scott Smith, and James Zou. 2024. What’s documented in AI? Systematic Analysis of 32K AI Model Cards. arXiv:2402.05160 [cs.SE]
- Liu et al. (2023) Yunpeng Liu, Kexin Li, Zhuotao Liu, Bihan Wen, Ke Xu, Weiqiang Wang, Wenbiao Zhao, and Qi Li. 2023. Provenance of Training without Training Data: Towards Privacy-Preserving DNN Model Ownership Verification. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). Association for Computing Machinery, New York, NY, USA, 1980–1990. https://doi.org/10.1145/3543507.3583198
@@
- Mei et al. (2022) Songzhu Mei, Cong Liu, Qinglin Wang, and Huayou Su. 2022. Model Provenance Management in MLOps Pipeline. In Proceedings of the 2022 8th International Conference on Computing and Data Engineering (Bangkok, Thailand) (ICCDE ’22). Association for Computing Machinery, New York, NY, USA, 45–50. https://doi.org/10.1145/3512850.3512861

### v2

@@ 7. Conclusion
The database community has responded to the “Big Model" revolution by proposing platforms like Agora (Traub et al., 2020), that manage data-related assets, including models, datasets, software, and compute resources in a coherent ecosystem. But the model-specific support in such an ecosystem must be expanded to include general methods for managing and finding models, support that can make such systems fully functional model lakes. It will be important that the methods generalize, irrespective of how many models we are trying to understand, what architectures the models use, how they are trained, or how we wish to query or search the models. We call on the database community to contribute to the vision of model lakes, supporting users to more easily find relevant models and to better understand those models. Our vision is for a fundamentally new platform that extends and integrates work on data/model attribution, data/model search, and data/model version management.
@@ Acknowledgements.
KP and RM were supported in part by NSF award numbers IIS-2107248, IIS-1956096, IIS-2325632, and KP and DB by a grant from Open Philanthropy.References
@@
- Achille et al. (2019) Alessandro Achille, Michael Lam, Rahul Tewari, Avinash Ravichandran, Subhransu Maji, Charless Fowlkes, Stefano Soatto, and Pietro Perona. 2019. Task2Vec: Task Embedding for Meta-Learning. In 2019 IEEE/CVF International Conference on Computer Vision (ICCV). 6429–6438. https://doi.org/10.1109/ICCV.2019.00653
@@
- Akhtar et al. (2024) Mubashara Akhtar, Omar Benjelloun, Costanza Conforti, Pieter Gijsbers, Joan Giner-Miguelez, Nitisha Jain, Michael Kuchnik, Quentin Lhoest, Pierre Marcenac, Manil Maskey, Peter Mattson, Luis Oala, Pierre Ruyssen, Rajat Shinde, Elena Simperl, Goeffry Thomas, Slava Tykhonov, Joaquin Vanschoren, Jos van der Velde, Steffen Vogler, and Carole-Jean Wu. 2024. Croissant: A Metadata Format for ML-Ready Datasets. In Proceedings of the Eighth Workshop on Data Management for End-to-End Machine Learning (DEEM ’24). Association for Computing Machinery, New York, NY, USA, 1–6. https://doi.org/10.1145/3650203.3663326
- Andy Zou (2023) Sarah Chen James Campbell Phillip Guo Richard Ren Alexander Pan Xuwang Yin Mantas Mazeika Ann-Kathrin Dombrowski Shashwat Goel Nathaniel Li Michael J. Byun Zifan Wang Alex Mallen Steven Basart Sanmi Koyejo Dawn Song Matt Fredrikson Zico Kolter Dan Hendrycks Andy Zou, Long Phan. 2023. Representation Engineering: A Top-Down Approach to AI Transparency. arXiv:cs.CL/2310.01405
- Anil et al. (2023) Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M. Dai, Anja Hauth, Katie Millican, David Silver, Slav Petrov, Melvin Johnson, Ioannis Antonoglou, Julian Schrittwieser, Amelia Glaese, Jilin Chen, Emily Pitler, Timothy P. Lillicrap, Angeliki Lazaridou, Orhan Firat, James Molloy, Michael Isard, Paul Ronald Barham, Tom Hennigan, Benjamin Lee, Fabio Viola, Malcolm Reynolds, Yuanzhong Xu, Ryan Doherty, Eli Collins, Clemens Meyer, Eliza Rutherford, Erica Moreira, Kareem Ayoub, Megha Goel, George Tucker, Enrique Piqueras, Maxim Krikun, Iain Barr, Nikolay Savinov, Ivo Danihelka, Becca Roelofs, Anaïs White, Anders Andreassen, Tamara von Glehn, Lakshman Yagati, Mehran Kazemi, Lucas Gonzalez, Misha Khalman, Jakub Sygnowski, and et al. 2023. Gemini: A Family of Highly Capable Multimodal Models. CoRR abs/2312.11805 (2023). https://doi.org/10.48550/ARXIV.2312.11805 arXiv:2312.11805
- Anthropic (2024) Anthropic. 2024. The Claude 3 Model Family: Opus, Sonnet, Haiku. https://www-cdn.anthropic.com/de8ba9b01c9ab7cbabf5c33b80b7bbc618857627/Model_Card_Claude_3.pdf. Accessed: 2024-09-30.
@@
- Belrose et al. (2023) Nora Belrose, David Schneider-Joseph, Shauli Ravfogel, Ryan Cotterell, Edward Raff, and Stella Biderman. 2023. LEACE: Perfect linear concept erasure in closed form. arXiv:cs.LG/2306.03819
- Bernstein and Melnik (2007) Philip A. Bernstein and Sergey Melnik. 2007. Model management 2.0: manipulating richer mappings. In Proceedings of the 2007 ACM SIGMOD International Conference on Management of Data (SIGMOD ’07). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/1247480.1247482
- Bogatu et al. (2020) Alex Bogatu, Alvaro A. A. Fernandes, Norman W. Paton, and Nikolaos Konstantinou. 2020. Dataset Discovery in Data Lakes. In ICDE. 709–720.
@@
- Buneman et al. (2001) Peter Buneman, Sanjeev Khanna, and Tan Wang-Chiew. 2001. Why and where: A characterization of data provenance. In Database Theory—ICDT 2001: 8th International Conference London, UK, January 4–6, 2001 Proceedings 8. Springer, 316–330.
@@
- Carlini et al. (2021) Nicholas Carlini, Florian Tramer, Eric Wallace, Matthew Jagielski, Ariel Herbert-Voss, Katherine Lee, Adam Roberts, Tom Brown, Dawn Song, Ulfar Erlingsson, et al. 2021. Extracting training data from large language models. In 30th USENIX Security Symposium (USENIX Security 21). 2633–2650.
- Casper et al. (2024) Stephen Casper, Carson Ezell, Charlotte Siegmann, Noam Kolt, Taylor Lynn Curtis, Benjamin Bucknall, Andreas Haupt, Kevin Wei, Jérémy Scheurer, Marius Hobbhahn, Lee Sharkey, Satyapriya Krishna, Marvin Von Hagen, Silas Alberti, Alan Chan, Qinyi Sun, Michael Gerovitch, David Bau, Max Tegmark, David Krueger, and Dylan Hadfield-Menell. 2024. Black-Box Access is Insufficient for Rigorous AI Audits. arXiv:cs.CY/2401.14446
@@
- Conmy et al. (2023) Arthur Conmy, Augustine N. Mavor-Parker, Aengus Lynch, Stefan Heimersheim, and Adrià Garriga-Alonso. 2023. Towards Automated Circuit Discovery for Mechanistic Interpretability. arXiv:cs.LG/2304.14997
- DeepSeek-AI et al. (2025) DeepSeek-AI, Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, Xiao Bi, Xiaokang Zhang, Xingkai Yu, Yu Wu, Z. F. Wu, Zhibin Gou, Zhihong Shao, Zhuoshu Li, Ziyi Gao, Aixin Liu, Bing Xue, Bingxuan Wang, Bochao Wu, Bei Feng, Chengda Lu, Chenggang Zhao, Chengqi Deng, Chenyu Zhang, Chong Ruan, Damai Dai, Deli Chen, Dongjie Ji, Erhang Li, Fangyun Lin, Fucong Dai, Fuli Luo, Guangbo Hao, Guanting Chen, Guowei Li, H. Zhang, Han Bao, Hanwei Xu, Haocheng Wang, Honghui Ding, Huajian Xin, Huazuo Gao, Hui Qu, Hui Li, Jianzhong Guo, Jiashi Li, Jiawei Wang, Jingchang Chen, Jingyang Yuan, Junjie Qiu, Junlong Li, J. L. Cai, Jiaqi Ni, Jian Liang, Jin Chen, Kai Dong, Kai Hu, Kaige Gao, Kang Guan, Kexin Huang, Kuai Yu, Lean Wang, Lecong Zhang, Liang Zhao, Litong Wang, Liyue Zhang, Lei Xu, Leyi Xia, Mingchuan Zhang, Minghua Zhang, Minghui Tang, Meng Li, Miaojun Wang, Mingming Li, Ning Tian, Panpan Huang, Peng Zhang, Qiancheng Wang, Qinyu Chen, Qiushi Du, Ruiqi Ge, Ruisong Zhang, Ruizhe Pan, Runji Wang, R. J. Chen, R. L. Jin, Ruyi Chen, Shanghao Lu, Shangyan Zhou, Shanhuang Chen, Shengfeng Ye, Shiyu Wang, Shuiping Yu, Shunfeng Zhou, Shuting Pan, S. S. Li, Shuang Zhou, Shaoqing Wu, Shengfeng Ye, Tao Yun, Tian Pei, Tianyu Sun, T. Wang, Wangding Zeng, Wanjia Zhao, Wen Liu, Wenfeng Liang, Wenjun Gao, Wenqin Yu, Wentao Zhang, W. L. Xiao, Wei An, Xiaodong Liu, Xiaohan Wang, Xiaokang Chen, Xiaotao Nie, Xin Cheng, Xin Liu, Xin Xie, Xingchao Liu, Xinyu Yang, Xinyuan Li, Xuecheng Su, Xuheng Lin, X. Q. Li, Xiangyue Jin, Xiaojin Shen, Xiaosha Chen, Xiaowen Sun, Xiaoxiang Wang, Xinnan Song, Xinyi Zhou, Xianzu Wang, Xinxia Shan, Y. K. Li, Y. Q. Wang, Y. X. Wei, Yang Zhang, Yanhong Xu, Yao Li, Yao Zhao, Yaofeng Sun, Yaohui Wang, Yi Yu, Yichao Zhang, Yifan Shi, Yiliang Xiong, Ying He, Yishi Piao, Yisong Wang, Yixuan Tan, Yiyang Ma, Yiyuan Liu, Yongqiang Guo, Yuan Ou, Yuduan Wang, Yue Gong, Yuheng Zou, Yujia He, Yunfan Xiong, Yuxiang Luo, Yuxiang You, Yuxuan Liu, Yuyang Zhou, Y. X. Zhu, Yanhong Xu, Yanping Huang, Yaohui Li, Yi Zheng, Yuchen Zhu, Yunxian Ma, Ying Tang, Yukun Zha, Yuting Yan, Z. Z. Ren, Zehui Ren, Zhangli Sha, Zhe Fu, Zhean Xu, Zhenda Xie, Zhengyan Zhang, Zhewen Hao, Zhicheng Ma, Zhigang Yan, Zhiyu Wu, Zihui Gu, Zijia Zhu, Zijun Liu, Zilin Li, Ziwei Xie, Ziyang Song, Zizheng Pan, Zhen Huang, Zhipeng Xu, Zhongyu Zhang, and Zhen Zhang. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:cs.CL/2501.12948 https://arxiv.org/abs/2501.12948
- Deng et al. (2023) Xiang Deng, Yu Gu, Boyuan Zheng, Shijie Chen, Samuel Stevens, Boshi Wang, Huan Sun, and Yu Su. 2023. Mind2Web: Towards a Generalist Agent for the Web. In Thirty-seventh Conference on Neural Information Processing Systems. https://openreview.net/forum?id=kiYqbO3wqw
@@
- Dibia (2023) Victor Dibia. 2023. LIDA: A Tool for Automatic Generation of Grammar-Agnostic Visualizations and Infographics using Large Language Models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations), Danushka Bollegala, Ruihong Huang, and Alan Ritter (Eds.). Association for Computational Linguistics, Toronto, Canada, 113–126. https://doi.org/10.18653/v1/2023.acl-demo.11
- Dong et al. (2022) Qingxiu Dong, Lei Li, Damai Dai, Ce Zheng, Zhiyong Wu, Baobao Chang, Xu Sun, Jingjing Xu, and Zhifang Sui. 2022. A Survey for In-context Learning. arXiv preprint arXiv:2301.00234 (2022).
@@
- Dou et al. (2024) Yao Dou, Isadora Krsek, Tarek Naous, Anubha Kabra, Sauvik Das, Alan Ritter, and Wei Xu. 2024. Reducing Privacy Risks in Online Self-Disclosures with Language Models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 13732–13754. https://doi.org/10.18653/v1/2024.acl-long.741
- Dwork et al. (2006) Cynthia Dwork, Frank McSherry, Kobbi Nissim, and Adam Smith. 2006. Calibrating noise to sensitivity in private data analysis. In Theory of Cryptography: Third Theory of Cryptography Conference, TCC 2006, New York, NY, USA, March 4-7, 2006. Proceedings 3. Springer, 265–284.
- Eilertsen et al. (2020) Gabriel Eilertsen, Daniel Jönsson, Timo Ropinski, Jonas Unger, and Anders Ynnerman. 2020. Classifying the classifier: dissecting the weight space of neural networks. In ECAI 2020. IOS Press, 1119–1126.
- Eldan and Russinovich (2023) Ronen Eldan and Mark Russinovich. 2023. Who’s Harry Potter? Approximate Unlearning in LLMs. arXiv:cs.CL/2310.02238
@@
- Fan et al. (2023a) Grace Fan, Jin Wang, Yuliang Li, and Renée J. Miller. 2023a. Table Discovery in Data Lakes: State-of-the-art and Future Directions. In Companion of the 2023 International Conference on Management of Data (SIGMOD ’23). Association for Computing Machinery, New York, NY, USA, 69–75. https://doi.org/10.1145/3555041.3589409
@@
- Golpayegani et al. (2024) Delaram Golpayegani, Isabelle Hupont, Cecilia Panigutti, Harshvardhan J Pandit, Sven Schade, Declan O’Sullivan, and Dave Lewis. 2024. AI cards: towards an applied framework for machine-readable AI and risk documentation inspired by the EU AI Act. In Annual Privacy Forum. Springer, 48–72.
@@
- Goyal et al. (2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh. 2017. Making the v in VQA Matter: Elevating the Role of Image Understanding in Visual Question Answering. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition (CVPR).
- Grosse et al. (2023) Roger Grosse, Juhan Bae, Cem Anil, Nelson Elhage, Alex Tamkin, Amirhossein Tajdini, Benoit Steiner, Dustin Li, Esin Durmus, Ethan Perez, et al. 2023. Studying large language model generalization with influence functions. arXiv preprint arXiv:2308.03296 (2023).
- Hendrycks et al. (2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2021a. Measuring Massive Multitask Language Understanding. In 9th International Conference on Learning Representations, ICLR 2021, Virtual Event, Austria, May 3-7, 2021. OpenReview.net. https://openreview.net/forum?id=d7KBjmI3GmQ
- Hendrycks et al. (2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. 2021b. Measuring Mathematical Problem Solving With the MATH Dataset. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2). https://openreview.net/forum?id=7Bywt2mQsCe
- Heusel et al. (2017) Martin Heusel, Hubert Ramsauer, Thomas Unterthiner, Bernhard Nessler, and Sepp Hochreiter. 2017. Gans trained by a two time-scale update rule converge to a local nash equilibrium. Advances in neural information processing systems 30 (2017).
- Horwitz et al. (2024) Eliahu Horwitz, Asaf Shul, and Yedid Hoshen. 2024. On the Origin of Llamas: Model Tree Heritage Recovery. arXiv preprint arXiv:2405.18432 (2024).
- House (2022) The White House. 2022. Blueprint for an AI Bill of Rights: Making Automated Systems Work for the American People. The White House.
@@
- Hu et al. (2023) Xuming Hu, Shen Wang, Xiao Qin, Chuan Lei, Zhengyuan Shen, Christos Faloutsos, Asterios Katsifodimos, George Karypis, Lijie Wen, and Philip S. Yu. 2023. Automatic Table Union Search with Tabular Representation Learning. In Findings of the Association for Computational Linguistics: ACL 2023, Toronto, Canada, July 9-14, 2023. Association for Computational Linguistics, 3786–3800. https://aclanthology.org/2023.findings-acl.233
- Huang et al. (2024) Xinting Huang, Madhur Panwar, Navin Goyal, and Michael Hahn. 2024. InversionView: A General-Purpose Method for Reading Information from Neural Activations. arXiv:cs.LG/2405.17653 https://arxiv.org/abs/2405.17653
@@
- Jiang et al. (2023) Albert Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de Las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, Lélio Renard Lavaud, Marie-Anne Lachaux, Pierre Stock, Teven Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed. 2023. Mistral 7B. CoRR abs/2310.06825 (2023). https://doi.org/10.48550/ARXIV.2310.06825 arXiv:2310.06825
@@
- Koh and Liang (2017) Pang Wei Koh and Percy Liang. 2017. Understanding black-box predictions via influence functions. In International conference on machine learning. PMLR, 1885–1894.
@@
- Labs ([n.d.]a) Neptune Labs. [n.d.]a. neptune.ai | The MLOps stack component for experiment tracking. https://neptune.ai/.
@@
- Lacoste et al. (2019) Alexandre Lacoste, Alexandra Luccioni, Victor Schmidt, and Thomas Dandres. 2019. Quantifying the Carbon Emissions of Machine Learning. arXiv preprint arXiv:1910.09700 (2019).
- Lenc and Vedaldi (2015) Karel Lenc and Andrea Vedaldi. 2015. Understanding image representations by measuring their equivariance and equivalence. In Proceedings of the IEEE conference on computer vision and pattern recognition. 991–999.
@@
- Li et al. (2025) Tianlong Li, Zhenghua Wang, Wenhao Liu, Muling Wu, Shihan Dou, Changze Lv, Xiaohua Wang, Xiaoqing Zheng, and Xuanjing Huang. 2025. Revisiting Jailbreaking for Large Language Models: A Representation Engineering Perspective. In Proceedings of the 31st International Conference on Computational Linguistics, Owen Rambow, Leo Wanner, Marianna Apidianaki, Hend Al-Khalifa, Barbara Di Eugenio, and Steven Schockaert (Eds.). Association for Computational Linguistics, Abu Dhabi, UAE, 3158–3178. https://aclanthology.org/2025.coling-main.212/
@@
- Liang et al. (2024) Weixin Liang, Nazneen Rajani, Xinyu Yang, Ezinwanne Ozoani, Eric Wu, Yiqun Chen, Daniel Scott Smith, and James Zou. 2024. What’s documented in AI? Systematic Analysis of 32K AI Model Cards. arXiv:cs.SE/2402.05160
- Lin et al. (2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick. 2014. Microsoft coco: Common objects in context. In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13. Springer, 740–755.
- Liu et al. (2023) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2023. Pre-train, Prompt, and Predict: A Systematic Survey of Prompting Methods in Natural Language Processing. ACM Comput. Surv. 55, 9, Article 195 (Jan. 2023), 35 pages. https://doi.org/10.1145/3560815
- Longpre et al. (2024) Shayne Longpre, Stella Biderman, Alon Albalak, Hailey Schoelkopf, Daniel McDuff, Sayash Kapoor, Kevin Klyman, Kyle Lo, Gabriel Ilharco, Nay San, Maribeth Rauh, Aviya Skowron, Bertie Vidgen, Laura Weidinger, Arvind Narayanan, Victor Sanh, David Ifeoluwa Adelani, Percy Liang, Rishi Bommasani, Peter Henderson, Sasha Luccioni, Yacine Jernite, and Luca Soldaini. 2024. The Responsible Foundation Model Development Cheatsheet: A Review of Tools & Resources. CoRR abs/2406.16746 (2024). https://doi.org/10.48550/ARXIV.2406.16746 arXiv:2406.16746
- Longpre et al. (2023) Shayne Longpre, Robert Mahari, Anthony Chen, Naana Obeng-Marnu, Damien Sileo, William Brannon, Niklas Muennighoff, Nathan Khazam, Jad Kabbara, Kartik Perisetla, Xinyi Wu, Enrico Shippole, Kurt D. Bollacker, Tongshuang Wu, Luis Villa, Sandy Pentland, Deb Roy, and Sara Hooker. 2023. The Data Provenance Initiative: A Large Scale Audit of Dataset Licensing & Attribution in AI. CoRR abs/2310.16787 (2023). https://doi.org/10.48550/ARXIV.2310.16787 arXiv:2310.16787
@@
- Malkov and Yashunin (2020) Yury A. Malkov and Dmitry A. Yashunin. 2020. Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs. IEEE Trans. Pattern Anal. Mach. Intell. 42, 4 (2020), 824–836.
- Mattson et al. (2019) Peter Mattson, Christine Cheng, Cody Coleman, Greg Diamos, Paulius Micikevicius, David A. Patterson, Hanlin Tang, Gu-Yeon Wei, Peter Bailis, Victor Bittorf, David Brooks, Dehao Chen, Debojyoti Dutta, Udit Gupta, Kim M. Hazelwood, Andrew Hock, Xinyuan Huang, Bill Jia, Daniel Kang, David Kanter, Naveen Kumar, Jeffery Liao, Guokai Ma, Deepak Narayanan, Tayo Oguntebi, Gennady Pekhimenko, Lillian Pentecost, Vijay Janapa Reddi, Taylor Robie, Tom St. John, Carole-Jean Wu, Lingjie Xu, Cliff Young, and Matei Zaharia. 2019. MLPerf Training Benchmark. CoRR abs/1910.01500 (2019). arXiv:1910.01500 http://arxiv.org/abs/1910.01500
- Mei et al. (2022) Songzhu Mei, Cong Liu, Qinglin Wang, and Huayou Su. 2022. Model Provenance Management in MLOps Pipeline. In Proceedings of the 2022 8th International Conference on Computing and Data Engineering (ICCDE ’22). Association for Computing Machinery, New York, NY, USA, 45–50. https://doi.org/10.1145/3512850.3512861